# HERACLES++: a multi-dimensional Eulerian code for exascale computing

### v1

@@ C.1 Sod shock tube test
The classical shock tube test is designed to demonstrate the code’s shock-capturing capabilities. We adopted 1-D Cartesian coordinates, an ideal gas with $\gamma=1.4$, and transmissive boundary conditions. The spatial domain is $x\in[0,1]$, sampled with 100 grid points. The initial conditions were taken from Toro (2009) and such that:
$(\rho,u,P)(x)=\left\{\begin{aligned} &(1.0,0,1.0)&\quad\textrm{if }x<0.5\\ &(0.125,0,0.1)&\quad\textrm{if }x>0.5\end{aligned}\right.$ (29)
This test is not particularly challenging for modern hydrodynamics codes but it is interesting because it exhibits all fundamental hydrodynamical waves. It is also well-documented. The exact solution is known analytically and consists of a propagating rarefaction wave on the left side, a rightward-propagating contact discontinuity, and a rightward-propagating shock.
@@
C.2 2-D Rayleigh-Taylor test
In Section 4.2, we presented results for the Rayleigh-Taylor instability test in 3-D. Here, we consider a similar problem but in 2-D and with different initial perturbations, which we took from Skinner et al. (2019).
The fluid is contained in a box of size defined by $x\in[-0.5,05]$, $z\in[-1,1]$, and with $n_{x}=1024$ and $n_{z}=2048$. The gravitational acceleration is constant with $g_{z}=-0.5$ and the initial density of the stationary fluid is given by:
@@
with $P_{0}=10/7+1/4$. For this test we adopted an ideal gas with $\gamma$ = 1.4. At time $t=0$, the interface between the two fluids was perturbed according to:
$h(x)=h_{0}\cos(\kappa x)$ (32)
@@
Figure 21 shows the increasing mixing between the two fluids as time progresses, as well as the formation of the characteristic mushroom structures typical of the Rayleigh-Taylor instability. With the heavier fluid above, the slightest disturbance of the interface amplifies exponentially. Fig. 22 compares the evolution of the size of the mixing layer according to the analytical theory: $h(t)=h_{0}\cosh{(\sqrt{Ag\kappa}\,t)}$, with $A=1/3$, the Atwood number. The result shows a great agreement with the linear theory up to tsimilar-$t\sim$1.75, when the Rayleigh-Taylor structures start to develop.
Figure 22: Amplitude of the interface perturbation for the single mode 2-D Rayleigh-Taylor test and comparison with the analytical solution.
@@ C.3 Liska-Wendroff implosion test
Figure 23: Results for the 2-D implosion test. We show a colormap of the density at $t=2.5$, with the $x=z$ symmetry axis in black dashed line. (see Section C.3 for details.)
The 2-D implosion test of Liska & Wendroff (2003) consists of Sod-like initial conditions, but with a 45 deg rotation. The boundary conditions are reflective in every direction and the size of the domain is $(x,z)\in[0,0.3]$, which we cover with 1024 zones in each direction. We again adopted an ideal gas with $\gamma=1.4$. Following Liska & Wendroff (2003), the initial conditions of the stationary fluid are:
@@ C.4 Multidimensional Sedov Blast wave
In Section 4.3, we present a 3-D version of the off-center Sedov blast wave in the spherical coordinate. We also test the Sedov explosion in spherical 1-D and 2-D, with a centered explosion. We compare with the analytical solution.
For the 1-D simulation, the domain is $r\in[0,1]$ (which we sample with 400 grid points) and the explosion is initiated at $r=0$. The initial conditions are $\rho=1,\quad u=0$ and $\gamma=1.4$. $E=E_{0}=10^{-12}$ everywhere except at the point of explosion where $E=E_{1}=1$ at the point of the explosion. The boundary conditions are reflexive at the center and transmissive at the maximum radius. The 2-D counterpart is started from the results of the 1-D simulation at $t=0.3$, with the fluid properties copied across all polar directions. The spherical grid is such that $r\in[0.3,1.3]$, $\theta\in[\pi/4,3\pi/4]$, with 400 grid points in each direction.
Figure 24 compares the density and total energy at $t=0.7$ for the 1-D and 2-D simulations (we also added the analytical solution for each quantity). The position of the shock predicted by HERACLES++ is similar in 1-D and 2-D and it also agrees with the analytical solution. Figure 25 shows the density at $t=0.7$. The shock remains spherical and its outer radius follows the analytical prediction (Eq. 25, with $\xi_{0}\approx 1.15$).
Figure 24: Comparison of the density (top) and total energy (bottom) at $t=0.7$ for the Sedov blast wave test performed with HERACLES++ in 1-D (blue) and 2-D (pink; we show the result along $\theta=\pi/2$). We also show the analytic solution (black dashed line).
Figure 25: Colormap of the density at $t=0.7$ for the 2-D Sedov blast wave test (the analytic solution is shown as a black line).

### v2

@@ C.1 Sod shock tube test
The classical shock tube test is designed to demonstrate the code’s shock-capturing capabilities. We adopted 1D Cartesian coordinates, an ideal gas with $\gamma=1.4$, and transmissive boundary conditions. The spatial domain is $x\in[0,1]$, sampled with 100 grid points. The initial conditions were taken from Toro (2009) and such that
$(\rho,u,P)(x)=\left\{\begin{aligned} &(1.0,0,1.0)&\quad\textrm{if }x<0.5\\ &(0.125,0,0.1)&\quad\textrm{if }x>0.5.\end{aligned}\right.$ (29)
This test is not particularly challenging for modern hydrodynamics codes but it is interesting because it exhibits all fundamental hydrodynamical waves. It is also well documented. The exact solution is known analytically and consists of a propagating rarefaction wave on the left side, a rightward-propagating contact discontinuity, and a rightward-propagating shock.
@@
C.2 2D Rayleigh-Taylor test
In Sect. 4.2 we presented results for the Rayleigh-Taylor instability test in 3D. Here, we consider a similar problem but in 2D and with different initial perturbations, which we took from Skinner et al. (2019).
The fluid is contained in a box of size defined by $x\in[-0.5,05]$, $z\in[-1,1]$, and with $n_{x}=1024$ and $n_{z}=2048$. The gravitational acceleration is constant with $g_{z}=-0.5$ and the initial density of the stationary fluid is given by
@@
with $P_{0}=10/7+1/4$. For this test we adopted an ideal gas with $\gamma$ = 1.4. At time $t=0$, the interface between the two fluids was perturbed according to
$h(x)=h_{0}\cos(\kappa x),$ (32)
@@
Figure 21 shows the increasing mixing between the two fluids as time progresses, as well as the formation of the characteristic mushroom structures typical of the Rayleigh-Taylor instability. With the heavier fluid above, the slightest disturbance of the interface amplifies exponentially. Figure 22 compares the evolution of the size of the mixing layer according to the analytical theory: $h(t)=h_{0}\cosh{(\sqrt{Ag\kappa}\,t)}$, with $A=1/3$, the Atwood number. The result shows a great agreement with the linear theory up to tsimilar-$t\sim$1.5, when the Rayleigh-Taylor structures start to develop.
Figure 22: Amplitude of the interface perturbation for the single mode 2D Rayleigh-Taylor test and comparison with the analytical solution.
@@ C.3 Liska-Wendroff implosion test
Figure 23: Results for the 2D implosion test. We show a color map of the density at $t=2.5$, with the $x=z$ symmetry axis in black dashed line. (see Appendix C.3 for details.)
The 2D implosion test of Liska & Wendroff (2003) consists of Sod-like initial conditions, but with a 45 deg rotation. The boundary conditions are reflective in every direction and the size of the domain is $(x,z)\in[0,0.3]$, which we cover with 1024 zones in each direction. We again adopted an ideal gas with $\gamma=1.4$. Following Liska & Wendroff (2003), the initial conditions of the stationary fluid are
@@ C.4 Multidimensional Sedov blast wave
In Sect. 4.3 we present a 3D version of the off-center Sedov blast wave in the spherical coordinate. We also test the Sedov explosion in spherical 1D and 2D, with a centered explosion. We compare with the analytical solution.
For the 1D simulation, the domain is $r\in[0,1]$ (which we sample with 400 grid points) and the explosion is initiated at $r=0$. The initial conditions are $\rho=1,\quad u=0$ and $\gamma=1.4$. $E=E_{0}=10^{-12}$ everywhere except at the point of explosion where $E=E_{1}=1$ at the point of the explosion. The boundary conditions are reflexive at the center and transmissive at the maximum radius. The 2D counterpart is started from the results of the 1D simulation at $t=0.3$, with the fluid properties copied across all polar directions. The spherical grid is such that $r\in[0.3,1.3]$, $\theta\in[\pi/4,3\pi/4]$, with 400 grid points in each direction.
Figure 24 compares the density and total energy at $t=0.7$ for the 1D and 2D simulations (we also added the analytical solution for each quantity). The position of the shock predicted by HERACLES++ is similar in 1D and 2D and it also agrees with the analytical solution. Figure 25 shows the density at $t=0.7$. The shock remains spherical and its outer radius follows the analytical prediction (Eq. 25, with $\xi_{0}\approx 1.15$).
Figure 24: Comparison of the density (top) and total energy (bottom) at $t=0.7$ for the Sedov blast wave test performed with HERACLES++ in 1D (blue) and 2D (pink; we show the result along $\theta=\pi/2$). We also show the analytic solution (black dashed line).
Figure 25: Color map of the density at $t=0.7$ for the 2D Sedov blast wave test (the analytic solution is shown as a black dashed line).